A speedboat starts from rest and accelerates at +2.01 m/s2 for 7.10 s. At the end of this time, the boat continues for an additional 6.15 s with an acceleration of +0.518 m/s2. Following this, the boat accelerates at -1.49 m/s2 for 8.70 s.

(a) What is the velocity of the boat at t = 21.95 s? _______ m/s

(b) Find the total displacement of the boat. _______ m
Answer #1

v1 = 2.01*7.1 = 14.271 m/s

v2 = 14.271+0.518*6.15 = 17.46 m/s

v3 = 17.46-1.49*8.7 = 4.497 m/s

b) s1 = 1/2a1t1^2 = 50.7 m

s2 = 14.271*6.15+1/2*0.518*6.15^2 = 87.8+9.79 = 97.59 m

s3 = 17.46*8.7-1/2*1.49*8.7^2 =151.9- 56.39 = 95.51 m

s = 243.8 m